OpenStudy (anonymous):

(Please Help!) (Medal Reward!) What is the height of the following triangular prism if the surface area is 70.4 sq. in.? Surface Area: 70.4 Length: 3.6 Width: 2 Height: h Formula: SA=2lw+2lh+2wh

OpenStudy (anonymous):

we need a calculator... lets work this out together... 70.4=2*3.6*2 + 2*3.6*h + 2*2*h Do you know how to simplify that? All I did was fill in what we know...

OpenStudy (anonymous):

70.4=14.4+7.2h+4h ?

OpenStudy (anonymous):

I don't have a calculator but that looks right to me... you can add the 7.2h to the 4h, so you have 70.4=14.4+11.2h. Now you want to subtract 14.4 from both sides... What do you have?